The width of a rectangle is x inches long. The length of that rectangle is 5 inches less than twice the width. If the perimeter measures 26 inches, what is the measure of the length and width of the rectangle. Show all the work and the alegbra you used to solve this problem.

I am having a hard time. I'm not even sure where to start with this.
If 26 inches is the whole measurement and the length is 5 inches less than the width... Then would it be worked as this .... 5-x=26 ??? That doesn't seem right because once you work the problem it would be
5-x=26
+5 +5
x=31
31 could not be the measure of the length because that is greater than 26, which is the whole measurement of the unit.
Help please.

Answer by sudhanshu_kmr(1152)


width = x
length = 2*x-5
perimeter = 2*(x + 2x-5)

According to question..
2*(x+2x-5) = 26
=> 6x -10 =26
=> x = 36/6 = 6
width = 6 inches
length = 7 inches.
if any doubt, feel free to contact ....